A flat belt drive is to be used to drive a reciprocating compressor running at 720 rpm by a 15 kw 1440 rpm electric motor the required centre distance is 2 m. select the flat belt for the drive from the manufacturer's catalogue
1
Expert's answer
2021-08-05T03:19:22-0400

N= 720 rpm, P= 15 kW, Electric motor = 1440 rpm, center distance is 2 m

We will use belt velocity as 18 m/s .

"v_1= \frac{\pi d_1N}{60}"


"18 = \frac{\pi d 1440}{60}"

d= "\frac{18\times 60 }{\pi \times 1440}=0.238 m= 238 mm"

Load correction factor is 1.3 for compressor from table

So,

Maximum power = "1.3 \times 15 =" 19.5 kW

Step 2: Arc of contact factor

"\alpha _s= 180 - 2sin^{-1}(\frac{(D-d)}{2C})"

"\alpha_s=175.23 ^o"

and the arc contact factor for this angle

Fd = 1.019

Step 3:

Corrected power = 1.019 "\times maximum [power" = 19.87 kW

Step 4:

Corrected belt rating = "\frac{0.0118 \times 18.85 }{5.08}=0.0438 kW"

Step 5: Selection of belt

width "\times" No of piles = corrected power/ correct belt rating


4 piles , w ="\frac{756.17}{4}=189.04 mm"


5 piles , w="\frac{756.17}{5}=151 .2 mm"

We will select a HI-SPeed belt of 152 mm preferred width with 5 piles.